OpenStudy (anonymous):

Given the function f(x) = 5x - 9, find f(-3).

OpenStudy (shaik0124):

f(x)=5(-3)-9 =-15-9 =-24

OpenStudy (waleed_imtiaz):

put x=-3..... f(-3)=5(-3)-9=-15-9=-24

OpenStudy (shaik0124):

i replied already@waleed

OpenStudy (waleed_imtiaz):

yes.... I saw after my reply